# Jinhang Chai     Jianqing Fan

### v1

@@ Abstract
The problem of structured matrix estimation has been studied mostly under strong noise dependence assumptions. This paper considers a general framework of noisy low-rank-plus-sparse matrix recovery, where the noise matrix may come from any joint distribution with arbitrary dependence across entries. We propose an incoherent-constrained least-square estimator and prove its tightness both in the sense of deterministic lower bound and matching minimax risks under various noise distributions. To attain this, we establish a novel result asserting that the difference between two arbitrary low-rank incoherent matrices must spread energy out across its entries, in other words cannot be too sparse, which sheds light on the structure of incoherent low-rank matrices and may be of independent interest. We then showcase the applications of our framework to several important statistical machine learning problems. In the problem of estimating a structured Markov transition kernel, the proposed method achieves the minimax optimality and the result can be extended to estimating the conditional mean operator, a crucial component in reinforcement learning. The applications to multitask regression and structured covariance estimation are also presented. We propose an alternating minimization algorithm to approximately solve the potentially hard optimization problem. Numerical results corroborate the effectiveness of our method which typically converges in a few steps.
Keywords and Phrases: Incoherent-constrained least-squares; deterministic bounds; minimax optimality; Markovian processes; frequency matrices; transition mean.
@@ 1 Introduction
Over the past decade, the structures of low-rank-plus-sparse matrices and their variants have received widespread attention in statistical machine learning and been applied to a wide spectrum of scientific problems (Chen et al., 2021a, ), including matrix denoising (Chandrasekaran et al.,, 2011; Agarwal et al.,, 2012), robust PCA (Candès et al.,, 2011; Wright et al.,, 2009), multitask regression (Yuan et al.,, 2007; Rohde & Tsybakov,, 2011), factor models (Fan et al.,, 2013; Fan et al., 2021a, ), among others. Broadly speaking, the general goal of this problem is to recover the statistical signal which consists of a low-rank matrix plus a sparse matrix from noisy observations. Nonetheless, prevailing works typically impose strong assumptions on the structure of noise. They either posit the noise matrix exhibits independence across its entries (Candes & Plan,, 2010; Wang & Fan,, 2022), or that each entry possesses either a light tail distribution or a sub-Gaussian distribution(Chen et al., 2021b, ). Some also assume the signal-to-noise ratio is lower bounded (Chen et al., 2021b, ).
@@
Though near-optimal rates can be achieved, they rely on strong noise assumptions which are likely to fail in the real world. To illustrate the limitations of the existing theory, let us take a detour to look at a particular problem that motivated our study. Specifically, we focus on Markov chain transition matrix estimation (Hao et al.,, 2018), under the assumption that the transition matrix admits a low-rank-plus-sparse structure. Statistically, the goal is to estimate the ground-truth transition matrix $P^{\star}\in\mathbb{R}^{p\times p}$ given a single trajectory of Markov chain $\{X_{0},X_{1},\cdots X_{n}\}$ [^2] Our theory can be readily extended to the case with multiple trajectories., under the assumption that the ground-truth transition matrix follows the low-rank-plus-sparse form $P^{\star}=L^{\star}+S^{\star}$, where $L^{\star}$ is a low-rank matrix and $S^{\star}$ is a sparse matrix with a small number of nonzero elements. This structure is motivated by the fact that unstructured $P^{\star}$ becomes notably difficult to estimate when the dimension $p$ is large (Wolfer & Kontorovich,, 2019, 2021) and exact low-rankness can be overly restrictive in many applications.
@@
The problem of dependence of noises across entries arises in solving the estimation problem. To see this, let us consider a slightly simpler problem first: estimating the frequency matrix $F^{\star}$. [^3] For technical issues, we will focus on recovering the frequency matrix $F^{\star}$, which is given by $\operatorname{diag}(\pi)P^{\star}$. The rationale behind using the frequency matrix is that it has a special form of empirical average which facilitates the application of concentration inequalities. On the other hand, there is a one-to-one correspondence between $F^{*}$ and $P^{*}$, and the results of estimating $P^{\star}$ can be readily deduced from the results of estimating $F^{\star}$. After constructing the empirical frequency matrix $\widetilde{F}$, the problem can be recast as estimating the $F^{\star}$ from its noisy version $\widetilde{F}=F^{\star}+W$ with the noise matrix $W=\widetilde{F}-F^{\star}$. One can readily see that the entries of $W$ are dependent, invalidating most existing works on matrix denoising.
@@
We briefly go through the main techniques in proving the deterministic bound. First of all, by virtue of the optimality condition, the Frobenius loss of the optimal solution can be bounded by the Frobenius loss of the ground truth. However, a cross term appears in further deduction which hinders a straightforward application of Cauchy’s inequality to get the desired bound. While Agarwal et al., (2012) directly bounds the cross term using the spikiness assumption which they impose, a trailing term emerges in their final result, leading to possible estimation inconsistency even when the sample size goes to infinity. Whether the consistent estimation in such a dependent noise setting remains unclear. This work answers positively this question via a different route. To fulfill the goal, we prove a pivotal separation lemma indicating that the difference between two incoherent low-rank matrices cannot be too sparse. In other words, incoherent low-rank matrices are in $L_{0}$-norm sense well-separated. Therefore, the cross-term can be bounded by a fraction of the squared estimation error in the Frobenius norm, which leads to the desired results.
@@
At the core of our results lies an insightful yet intuitive lemma, Lemma B.1, whose proof is quite technically involved. We divide the entire proof into several parts for clarity. The proof proceeds from simple and special cases to more general ones. Specifically, we first prove the case with equal singular values in Lemma B.5 and then generalize it to allow for different singular values, by utilizing techniques from linear programming. The result in the lemma is also optimal, in the sense that the minimum sparsity level of arbitrary difference of two incoherent matrices cannot be higher up to a constant, as illustrated in Proposition B.7.
@@
Subsequently, we apply this overarching theory to several problem instances. For the Markov transition matrix estimation problem, we instantiate our deterministic bound to get a probabilistic bound which provably attains the minimax lower bound (Zhang & Wang,, 2019). We further employed the framework to address other types of problems, encompassing multitask learning with low-rank-plus-sparse structure and robust structured covariance estimation.
@@
To complement the acquired upper bounds, we turn to lower bounds. Specifically, in our deterministic result, we argue the rates of all terms are of optimal order by proving a matching deterministic lower bound defined in Section 3.1. In various noise settings including the Markov transition noise setting, we establish minimax risk, thereby showcasing the results’ tightness.
@@
To summarize, our contributions are three-hold. First of all, we formulate a general low-rank-plus-sparse matrix denoising problem that encompasses most statistical problems involving matrix estimation where the noise distribution across entries can be arbitrarily dependent and distributed. We develop an incoherent-constrained least-square optimization method that provably attains minimax risk in various scenarios and the method can be implemented in practice using an alternating minimization algorithm. Secondly, we find that the difference between two arbitrary low-rank incoherent matrices has energy spread across entries, namely, cannot be too sparse. This intermediate result sheds light on the geometric structure of low-rank incoherent matrix space as a subset of the entire matrix space. Thirdly, we employ our general theory to solve an important problem of estimating Markov transition matrix. We prove that our method achieves minimax lower bound. Further, we extend the study to estimating conditional mean which is an essential part in reinforcement learning. We prove that this structure promotes the statistical rate if the value vector is random and independent of the Markov chain data. In addition, we apply our result to multitask regression and structured covariance estimation. Extension of our general theory to rectangular matrix estimand is presented in Appendix D.
@@
Most of the matrix completion or Robust PCA literature hinges on strong assumptions on the error matrix such as independence across entries or sub-Gaussianity. It would be interesting to derive similar deterministic results in a partially observable setting, and we leave that to future work.
@@ 1.1 Notation
Before proceeding, we clarify some notations used in this paper. For an operator $\mathfrak{X}$, denote $\mathfrak{X}^{\star}$ as its conjugate operator. Denote $\mathbb{R}^{p\times p}$ as the set of $p\times p$ matrices whose elements are real numbers. Denote $\mathbb{R}^{p\times p}_{\operatorname{diag}}$ as the diagonal matrix space in $\mathbb{R}^{p\times p}$. Denote ${\cal O}^{p\times p}$ as the orthogonal matrix space in $\mathbb{R}^{p\times p}$. We use $\|\cdot\|,\|\cdot\|_{F},\|\cdot\|_{\max},\|\cdot\|_{1},\|\cdot\|_{*},\|\cdot\|_{0}$ to denote the spectral norm, Frobenius norm, elementwise $\ell_{\infty}$-norm, elementwise $\ell_{1}$-norm, nuclear norm and number of nonzero elements, respectively. For two matrices $F,G\in\mathbb{R}^{p\times q}$, denote $\langle F,G\rangle=\operatorname{Tr}(FG^{\top})$ as the matrix inner product in Euclidean distance. For every $p\geq r$, denote $\mathcal{O}_{p,r}$ as the set of $p\times r$ matrices whose columns are normalized and orthogonal, further denote $\mathcal{O}_{p,r}^{\mu}$ as all $\mu$-incoherent matrices in $\mathcal{O}_{p,r}$. By convention, vectors are usually column vectors, except that discrete probability distributions are row vectors. We use bold case $\mathbf{1}_{p}$ to denote the vector in $\mathbb{R}^{p}$ with all elements equal to 1 and $e_{i}$ to represent the $i^{th}$ standard basis vector whose elements are 0 except the $i^{th}$ coordinate is 1. For two vectors or matrices $u,v$, $u\geq v$ means for each entry $u_{i}\geq v_{i}$. We write $f(n)\lesssim g(n)$ or $f(n)=\operatorname{O}(g(n))$ if there exists a constant $C$ such that $f(n)\leq Cg(n)$ for all $n\geq n_{0}$. Similarly, write $f(n)\gtrsim g(n)$ if there exists a constant $C$ such that $f(n)\geq Cg(n)$ for all $n\geq n_{0}$. $f(n)\asymp g(n)$ denotes that both $f(n)\lesssim g(n)$ and $f(n)\gtrsim g(n)$. $\mathcal{U}[0,1]$ stands for uniform distribution over $[0,1]$. We denote $\{1,2,\cdots,p\}$ as $[p]$. Constants $C,c,c_{0},c_{1}$ can change from line to line.
@@ 1.2 Outline
Section 2 discusses some related literature and Section 4.1 briefly review some basic materials on basic Markov chains. In Section 3, we present the problem setup and our method to solve general low-rank-plus-sparse matrix estimation. Section 3.1 is devoted to the results of our proposed estimator, Section 4.2 applies this method to the Markov transition estimation problem, and Section 4.3 extends the results of Markov transition estimation to conditional mean estimation. In Section 5.1 and 5.2, we list two further applications on multi task regression and structured covariance estimation. In Section 3.2, we consider a practical algorithm to solve the optimization problem. Section 6 presents the numerical results. Section 7 concludes the paper. We present the proofs of main results in Appendix A, then elaborate on the proof of the key separation lemma in Appendix B. We include in Appendix C the proofs in Section 5. Appendix D is devoted to extension of rectangular estimand. We relegate the technical lemmas in the Appendix E.
2 Related work
Our work is related to the following lines of research directions.
  • Noisy Matrix Recovery: Noisy matrix recovery has been studied intensively in different settings such as noisy matrix completion (Candes & Plan,, 2010; Candes & Recht,, 2012; Gross,, 2011; Chen et al.,, 2019; Chen et al., 2020b, ) and robust PCA (Candès et al.,, 2011; Chandrasekaran et al.,, 2011; Chen et al., 2021b, ). There are mainly two popular ways to formulate the estimation problem, namely convex relaxation (Negahban & Wainwright,, 2012; Koltchinskii et al.,, 2011; Chen et al., 2020b, ; Chen et al., 2021b, ) and nonconvex Burer–Monteiro approach with gradient descent(Sun & Luo,, 2016; Ma et al.,, 2018; Chen et al., 2020a, ) or projected descent(Chen & Wainwright,, 2015; Zheng & Lafferty,, 2016). See overview papers (Chi et al.,, 2019; Chen & Chi,, 2018). With no missing entries, low-rank-and-sparse noisy matrix recovery has been studied in Hsu et al., (2011); Agarwal et al., (2012); Chandrasekaran et al., (2011); Netrapalli et al., (2014). Among those, one of the most relevant works to us is Agarwal et al., (2012), where the problem of study is under the name of noisy matrix decomposition. Other related matrix or vector sensing problems include: blind deconvolution(Ma et al.,, 2018; Chen et al.,, 2023), phase retrival(Ma et al.,, 2018), tensor decomposition(Ge & Ma,, 2017), etc.
  • Markov Transition Estimation: There has been a flurry of research on transition estimation. Some works consider the plain tabular case with no statistical structures imposed, such as Hao et al., (2018), Wolfer & Kontorovich, (2019), and Wolfer & Kontorovich, (2021). However, the rate inevitably depends on the size of the state space, which can be huge. To mitigate this dimension issue, Zhang & Wang, (2019) imposes low-rankness and derive the minimax rate of estimation, see also Zhu et al., (2022). To the best of our knowledge, our work is the first to consider the low-rank-plus-sparse structure on the transition matrix.
  • Spectral methods: Spectral methods have become a paradigm for many statistical machine problems (Chen et al., 2021a, ), and been a go-to choice for various kinds of problems, including PCA and factor analysis (Johnstone & Paul,, 2018; Fan et al.,, 2018), community detection (Abbe,, 2017), clustering (Ndaoud,, 2018; Löffler et al.,, 2021), ranking(Chen & Suh,, 2015), covariance estimation (Fan et al.,, 2013), etc. Although our approach does not utilize eigenvector perturbation theory due to the existence of a potentially large sparse part, the techniques bear a lot of similarity with the spectral methods which intensively use singular value decomposition and matrix analysis.
Here we give a detailed comparison of a few more related literature. Agarwal et al., (2012) bears the most resemblance to our work, which studied under a similar setting. Notably, their intermediate deterministic bound also accounts for arbitrary noise. However, they imposed a spikiness condition that each entry of the low-rank part of the ground-truth matrix is upper bounded in magnitude and the final deterministic bound involves a trailing term as appeared in Table 1. A huge drawback is that this bound results in inconsistency where the estimation error is lower bounded from zero even with infinite samples. Our result improves this.
Hsu et al., (2011) also considered the arbitrary noisy setting and gave bounds on estimating errors. However, compared to our work, there are several differences. Firstly, their bound is not tight in general. In particular, their bound on low-rank part is not optimal in their Theorem 3 and they can’t get optimal Frobenius norm bound while ours can. Secondly, our results apply to every linear observation model with restricted strong convexity while their results only apply to the standard low-rank-plus-sparse matrix estimation. Thirdly, unlike their work, we don’t need prior knowledge about error scale. We have less hyperparameters to tune than theirs. Again, our conditions on the sparse part are more relaxed than theirs.
There is a more recent work on Robust PCA (Chen et al., 2021b, ) which considered the setting where noise is i.i.d. and sub-Gaussian across entries. As previously mentioned, our results don’t need such an assumption. Their techniques of analyzing the rate convergence of gradient descent using leave-one-out highly rely on the assumption of independence of noise across entries. Moreover, unlike theirs, we can discard the assumption on the randomness of the sparse locations which is considered as fixed in our setting.
In constrast to aforementioned papers who considered noisy observations, Chandrasekaran et al., (2011) and Candès et al., (2011) studied under the noiseless setting where they proved with high probability that exact recovery is possible. In contrast, our setting involves noise and therefore fully recovering the true signal is impossible. Our conditions on the sparse part are more stringent than theirs.
Note that the key separation lemma B.1 characterizes the sparsity of arbitrary difference of two incoherent low-rank matrices. One can impose a more fine-grained structure on sparsity such as row-wise or column-wise sparsity (Chandrasekaran et al.,, 2011) and prove similar separation results. This is also an interesting future direction.

### v2

@@ Abstract
The problem of structured matrix estimation has been studied mostly under strong noise dependence assumptions. This paper considers a general framework of noisy low-rank-plus-sparse matrix recovery, where the noise matrix may come from any joint distribution with arbitrary dependence across entries. We propose an incoherent-constrained least-square estimator and prove its tightness both in the sense of deterministic lower bound and matching minimax risks under various noise distributions. To attain this, we establish a novel result asserting that the difference between two arbitrary low-rank incoherent matrices must spread energy out across its entries; in other words, it cannot be too sparse, which sheds light on the structure of incoherent low-rank matrices and may be of independent interest. We then showcase the applications of our framework to several important statistical machine learning problems. In the problem of estimating a structured Markov transition kernel, the proposed method achieves the minimax optimality and the result can be extended to estimating the conditional mean operator, a crucial component in reinforcement learning. The applications to multitask regression and structured covariance estimation are also presented. We propose an alternating minimization algorithm to approximately solve the potentially hard optimization problem. Numerical results corroborate the effectiveness of our method which typically converges in a few steps.
@@ 1 Introduction
Over the past decade, the structures of low-rank-plus-sparse matrices and their variants have received widespread attention in statistical machine learning and been applied to a wide spectrum of scientific problems (Chen et al., 2021a, ), including matrix denoising (Chandrasekaran et al.,, 2011; Agarwal et al.,, 2012), robust PCA (Candès et al.,, 2011; Wright et al.,, 2009), multitask regression (Yuan et al.,, 2007; Rohde & Tsybakov,, 2011), factor models (Fan et al.,, 2013; Fan et al., 2021a, ), among others. Broadly speaking, the general goal of this problem is to recover the statistical signal, which consists of a low-rank matrix plus a sparse matrix from noisy observations. Nonetheless, prevailing works typically impose strong assumptions on the structure of noise. They either posit the noise matrix exhibits independence across its entries (Candes & Plan,, 2010; Wang & Fan,, 2022), or that each entry possesses either a light tail distribution or a sub-Gaussian distribution (Chen et al., 2021b, ). Some also assume the signal-to-noise ratio is lower bounded (Chen et al., 2021b, ).
@@
Though near-optimal rates can be achieved, they rely on strong noise assumptions, which are likely to fail in the real world. To illustrate the limitations of the existing theory, let us take a detour to look at a particular problem that motivated our study. Specifically, we focus on Markov chain transition matrix estimation (Hao et al.,, 2018), under the assumption that the transition matrix admits a low-rank-plus-sparse structure. Statistically, the goal is to estimate the ground-truth transition matrix $P^{\star}\in\mathbb{R}^{p\times p}$ given a single trajectory of Markov chain $\{X_{0},X_{1},\cdots X_{n}\}$ [^2] Our theory can be readily extended to the case with multiple trajectories., under the assumption that the ground-truth transition matrix follows the low-rank-plus-sparse form $P^{\star}=L^{\star}+S^{\star}$, where $L^{\star}$ is a low-rank matrix and $S^{\star}$ is a sparse matrix with a small number of nonzero elements. This structure is motivated by the fact that unstructured $P^{\star}$ becomes notably difficult to estimate when the dimension $p$ is large (Wolfer & Kontorovich,, 2019, 2021) and exact low-rankness can be overly restrictive in many applications.
@@
The problem of dependence of noises across entries arises in solving the estimation problem. To see this, let us consider a slightly simpler problem first: estimating the frequency matrix $F^{\star}$. [^3] For technical issues, we will focus on recovering the frequency matrix $F^{\star}$, whose $(i,j)$ entry is $P(X_{t}=i,X_{t+1}=j)$ and is given by $\operatorname{diag}(\pi)P^{\star}$ with $\pi$ being the invariant distribution of the Markov chain. The rationale behind using the frequency matrix is that it has a special form of empirical average, which facilitates the application of concentration inequalities. On the other hand, there is an approximate one-to-one correspondence between the enlarged frequency matrix class (defined in (9)) and the transition matrix class, and the results of estimating $P^{\star}$ can be readily deduced from the results of estimating $F^{\star}$ in the enlarged frequency matrix class. After constructing the empirical frequency matrix $\widetilde{F}$, the problem can be recast as estimating the $F^{\star}$ from its noisy version $\widetilde{F}=F^{\star}+W$ with the noise matrix $W=\widetilde{F}-F^{\star}$. One can readily see that the entries of $W$ are dependent, invalidating most existing works on matrix denoising.
@@
We briefly go through the main techniques in proving the deterministic bound. First of all, by virtue of the optimality condition, the Frobenius loss of the optimal solution can be bounded by the Frobenius loss of the ground truth. However, a cross term appears in further deduction, which hinders a straightforward application of Cauchy-Schwarz inequality to get the desired bound. While Agarwal et al., (2012) directly bounds the cross term using the spikiness assumption which they impose, a trailing term emerges in their final result, leading to possible estimation inconsistency even when the sample size goes to infinity. It remains unclear whether consistent estimation is achievable in such a dependent noise setting. This work positively answers this question via a different route. To fulfill the goal, we prove a pivotal separation lemma indicating that the difference between two incoherent low-rank matrices cannot be too sparse. In other words, incoherent low-rank matrices are in $L_{0}$-norm sense well-separated. Therefore, the cross-term can be bounded by a fraction of the squared estimation error in the Frobenius norm, which leads to the desired results.
@@
At the core of our results lies an insightful yet intuitive Lemma D.1 whose proof is quite technically involved. We divide the entire proof into several parts for clarity. The proof proceeds from simple and special cases to more general ones. Specifically, we first prove the case with equal singular values in Lemma C.5 in the supplementary material and then generalize it to allow for different singular values by utilizing techniques from linear programming. The result in the lemma is also optimal, in the sense that the minimum sparsity level of arbitrary difference of two incoherent matrices cannot be higher up to a constant, as illustrated in Proposition C.7 in the supplementary material.
@@
Subsequently, we apply this overarching theory to several problem instances. For the Markov transition matrix estimation problem, we instantiate our deterministic bound to get a probabilistic bound, which provably attains the minimax lower bound (Zhang & Wang,, 2019). We further employed the framework to address other types of problems, encompassing multitask learning with low-rank-plus-sparse structure and robust structured covariance estimation.
@@
To complement the acquired upper bounds, we turn to lower bounds. Specifically, in our deterministic result, we argue the rates of all terms are of optimal order by proving a matching deterministic lower bound. In various noise settings including the Markov transition noise setting, we establish minimax risk, thereby showcasing the results’ tightness.
@@
To summarize, our contributions are three-fold. First of all, we formulate a general low-rank-plus-sparse matrix denoising problem that encompasses most statistical problems involving matrix estimation where the noise distribution across entries can be arbitrarily dependent and distributed. We develop an incoherent-constrained least-square optimization method that provably attains minimax risk in various scenarios, and the method can be implemented in practice using an alternating minimization algorithm. Secondly, we find that the difference between two arbitrary low-rank incoherent matrices has energy spread across entries, namely, cannot be too sparse. This intermediate result sheds light on the geometric structure of low-rank incoherent matrix space as a subset of the entire matrix space. Thirdly, we employ our general theory to solve an important problem of estimating the Markov transition matrix. We prove that our method achieves the minimax lower bound. Further, we extend the study to the estimation of the conditional mean, which is an essential part in reinforcement learning. We prove that this structure promotes the statistical rate if the value vector is random and independent of the Markov chain data. In addition, we apply our result to multitask regression and structured covariance estimation. An extension of our general theory to rectangular matrix estimand is presented in Appendix F in the supplementary material.
@@
Most of the matrix completion or Robust PCA literature hinges on strong assumptions on the error matrix, such as independence across entries or sub-Gaussianity. It would be interesting to derive similar deterministic results in a partially observable setting, and we leave that to future work.
@@ 1.1 Notation
Before proceeding, we clarify some notations used in this paper. For an operator $\mathfrak{X}$, denote $\mathfrak{X}^{*}$ as its conjugate operator. Denote $\mathbb{R}^{p\times p}$ as the set of $p\times p$ matrices whose elements are real numbers. Denote $\mathbb{R}^{p\times p}_{\operatorname{diag}}$ as the diagonal matrix space in $\mathbb{R}^{p\times p}$. Denote ${\cal O}^{p\times p}$ as the orthogonal matrix space in $\mathbb{R}^{p\times p}$. We use $\|\cdot\|,\|\cdot\|_{F},\|\cdot\|_{\max},\|\cdot\|_{1},\|\cdot\|_{*},\|\cdot\|_{0}$ to denote the spectral norm, Frobenius norm, elementwise $\ell_{\infty}$-norm, elementwise $\ell_{1}$-norm, nuclear norm and number of nonzero elements, respectively. We also denote $\|\cdot\|_{2,\infty}$ denote the $2$-to-$\infty$ norm, i.e., the maximum of the 2-norm of rows. For two matrices $F,G\in\mathbb{R}^{p\times q}$, denote $\langle F,G\rangle=\operatorname{Tr}(FG^{\top})$ as the matrix inner product in Euclidean distance. For every $p\geq r$, denote $\mathcal{O}_{p,r}$ as the set of $p\times r$ matrices whose columns are normalized and orthogonal, and $\mathcal{O}_{p,r}^{\mu}$ as in Definition 2.3. By convention, vectors are usually column vectors, except that discrete probability distributions are row vectors. We use bold case $\mathbf{1}_{p}$ to denote the vector in $\mathbb{R}^{p}$ with all elements equal to 1 and $e_{i}$ to represent the $i^{th}$ standard basis vector whose elements are 0 except the $i^{th}$ coordinate is 1. For two vectors or matrices $u,v$, $u\geq v$ means for each entry $u_{i}\geq v_{i}$. We write $f(n)\lesssim g(n)$ or $f(n)=\operatorname{O}(g(n))$ if there exists a constant $C$ such that $f(n)\leq Cg(n)$ for all $n\geq n_{0}$. Similarly, write $f(n)\gtrsim g(n)$ or $f(n)=\Omega(g(n))$ if there exists a constant $C$ such that $f(n)\geq Cg(n)$ for all $n\geq n_{0}$. $f(n)\asymp g(n)$ or $f(n)=\Theta(g(n))$ denotes that both $f(n)\lesssim g(n)$ and $f(n)\gtrsim g(n)$. We write $f(n,p)=\operatorname{\tilde{O}_{p}}(g(n,p))$ if there exists some $h(n,p)$ depending on logarithmic factors such that $\lim_{n\rightarrow\infty}\lim_{p\rightarrow\infty}\mathbb{P}(f(n,p)\leq g(n,p)h(n,p))=0$ We sometimes neglect logarithmic terms when specifying in the context. $\mathcal{U}[0,1]$ stands for uniform distribution over $[0,1]$. We denote $\{1,2,\cdots,p\}$ as $[p]$. We use ${\cal I}_{{\cal A}}$ or ${\cal I}({\cal A})$ to denote the indicator function of event ${\cal A}$. Constants $C,c,c_{0},c_{1}$ can change from line to line.
@@ 1.2 Outline
In Section 2, we present the problem setup and the method to solve general low-rank-plus-sparse matrix estimation. Section 3 focuses on the application to the estimation of frequency and transition matrix estimation. Section 4 illustrates further the versality of our method using the multi-task regression and robust stuctured matrix estimation. Section 5 presents the numerical results. Section 6 concludes the paper. The proofs and related works are delegated to the supplementary material.